A 1.0000 g sample that contains both sodium carbonate and sodium bicarbonate is dissolved in water and titrated with 0.1000 MCan you please solve this?
Answer #1

(a)

NaHCO3 (aq) -------> Na+ (aq) + HCO3-(aq)

Reaction and expression for Ka :

   HCO3-(aq) + H2O (l)\rightleftharpoons   H3O+ (aq) + CO32-(aq)

since water is in excess :

Ka = (H3O+)( CO32-)/( HCO3-)

Reaction and expression for Kb :

   HCO3-(aq) + H2O (l)\rightleftharpoons OH-(aq) + H2CO3(aq)

since water is in excess :

Kb = (OH-)( H2CO3)/( HCO3-)

(b) pH will be governed by various factors at various stage of titration :

1. pH at start of titration : (solution will be basic : pH : ~11)

principal equilibrium :

CO32-(aq) + H2O (l)\rightleftharpoons OH-(aq) + HCO3-(aq) ;   Kb = (OH-)( HCO3-)/( CO32-)

pH will be controlled by Kb and ( CO32-) .

2.before first equivalence point :

principal equilibrium :

CO32-(aq) + H3O+ (aq)\rightleftharpoons HCO3-(aq) + H2O (l)

It will behave as buffer of ( CO32-) and( HCO3-) : equation for pH of buffer will be used (pKa2).

3.pH at 1st equivalence point :(phenolphthalein end point )

pH = 1/2 (pKa1 + pKa2) = 8.3

4. between 1st and 2nd equivalence point :

principal equilibrium :

HCO3-(aq) + H3O+ (aq)\rightleftharpoons H2CO3(aq) + H2O (l)

equation for pH of buffer will be used (pKa1).

5. 2nd equivalence point : excess of HCl will determine pH.

(c) reaction during titration :

CO32-(aq) + H3O+ (aq)\rightleftharpoons HCO3-(aq) + H2O (l) (phenolphthalein end point : V1)

HCO3-(aq) + H3O+ (aq)\rightleftharpoons H2CO3(aq) + H2O (l) (methyl orange end point : V)

1.First point is solely due to Na2CO3 , so volume of HCl used for reacting with it : 2*V1

= 2*14.75 ml = 29.5 ml

moles of HCl reacted = 0.1000 mol/L * 29.5 ml / 1000 ml/L = 0.00295 moles

1 mol Na2CO3 = 2 mol of HCl

moles of Na2CO3 present = 0.001475 mole

mass of Na2CO3 present in sample = 0.001475 mole * 106 g/mol = 0.1564 g

2.

1. Volume of HCl reacted with NaHCO3 : V -2V1 = 9.35 ml

moles of HCl reacted = 0.1000 mol/L * 9.35 ml / 1000 ml/L = 0.000935 moles

1 mol NaHCO3 = 1 mol of HCl

moles of Na2CO3 present = 0.000935 moles

mass of NaHCO3 present in sample = 0.000935 moles * 84 g/mol = 0.0786 g